You have won the lottery. You will receive $2,550,000 today, and then receive 40 payments of $1,275,000 These payments will start one year from now and will be paid every six months. A representative from Greenleaf Investments has offered to purchase all the payments from you for $20 million. The appropriate discount rate is an APR of 10 percent compounded daily. Assume there are 12 months in a year, each with 30 days. What is the present value of the cash flows from your lottery winnings?
Answer #1

PV of 40 payments of $1275000,

Formula = P*e^tr, t= 40 payments & r=10% / 2 = 5% or 0.05 (being 6 months payments)

PV of 40 payments of $1275000= 1275000 * e^(40*0.05) = 1275000 * 7.3890561 = 9421046

Total PV of cash flows from the lottery winnings = today's payment + PV of 40 payments of $1275000

= 2550000 + 9421046

= $11,971,046